# Quantum Circuit Reconstruction from Power Side-Channel Attacks on Quantum Computer Controllers

### v1

@@ 1 Introduction
Power side channels are well studied for classical computers [EPMS23, PR13, WSRW21, ${}^{+}$21, ${}^{+}$20, ABP19, WSW19, BCHC18, BYT17, ${}^{+}$13, EW14, MOPT12]. There are even platforms [OC14] for analyzing power side channels. However, understanding power side channels for quantum computers has not been explored yet, which this work aims to address. We show how a malicious attacker can reconstruct a secret quantum circuit that is being executed, by simply measuring the power consumption.
@@ 1.1 Power Side-Channel Threats to Quantum Computers
The vulnerabilities in quantum computer controllers encompass more than just gate recovery. An attacker might discern the number of qubits used in a quantum program, a significant concern for algorithms like Quantum Approximate Optimization Algorithm (QAOA) [CK19]. Variational quantum algorithms [${}^{+}$21] are notably sensitive to qubit count and circuit depth. Extracting such hyperparameters can reveal crucial information about the algorithm. Moreover, given the current state of quantum computing where inputs are hard-coded, an attacker could potentially extract sensitive input data from the circuit.
@@ 1.2 Lessons from Historical Technological Threats
The significance of vulnerabilities in quantum computer controllers cannot be overstated, considering the ongoing evolution of technological advancements and their inherent risks. For perspective, speculative execution attacks in classical computers were not recognized as threats until 2018 [${}^{+}$20, ${}^{+}$20], despite having been operational since the commercialization of the technology by IBM and Intel in the 1990s. This delay between innovation and the identification of vulnerabilities underscores the necessity for proactive security measures.
@@ 3.1 Assumptions of Attacker Measurement
A weaker attacker could collect a single total power trace over all channels, but not distinguish the power traces of each channel. Such attackers can attempt Circuit Reconstruction attack by collecting a single total power trace (see Figure 7). In particular, there is a trend to have multiple AWGs being part of same physical device. For example, in QICK [${}^{+}$22] framework, FPGAs are used for waveform generation and one FPGA can generate many control pulses. We believe that going forward total power side-channel attacks may be most realistic as attacker may not easily get power traces for individual channel generated by the FPGA, but can easily measure total power consumption of the FPGA, and thus get total power trace of all the channels. Nevertheless, we explore both per-channel and total power side-channel attacks to understand their potential threats.
@@ 5.1 Single-shot Per-Channel Power Side-Channel Attack
where $v_{i}(x)$ and $Total_{A_{P}}(x)_{i}$ are the $i$-th elements of $v(x)$ and $Total_{A_{P}}(x)$ respectively. However, we evaluated various distances and metrics on randomly generated circuits including Euclidean distance and found that the Jensen-Shannon distance is the most suitable one for quantifying the candidate instructions against the measurements. This distance metric is commonly used to measure the dissimilarity between probability distributions and is well-suited for the task of comparing the probability distributions of pulse sequences. It has been applied to genome comparison[SJWK09, IHS10] in protein surface comparison [OR03], in machine learning [${}^{+}$20] and particularly in the analysis of the similarity between two quantum states [OBL22]. To compute the Jensen-Shannon distance, we first convert the trace and pulse data of a candidate instruction to two probability vectors $P$ and $Q$ by normalizing the amplitudes to turn them into discrete probability distributions. We then calculate the distance between the two probability vectors by computing the Jensen-Shannon distance (metric) between two probability arrays. This is the square root of the Jensen-Shannon divergence ($\sqrt{{\rm JSD}}$[ES03].
@@
In per-channel single trace attacks, we use the Kullback-Leibler divergence (KLD) and the Jensen-Shannon divergence (JSD) and its distance ($\sqrt{\text{JSD}}$) to distinguish two probability distributions. Therefore, here we provide a brief introduction to divergences, distances and metrics. For more details, we refer the reader to [DDDD09, ${}^{+}$14]. A metric $d$ on a set $\chi$ is a function $d:\chi\times\chi\rightarrow\mathbb{R}_{\geq 0}$ such that for any $x,y,z\in\chi$ the following properties are satisfied:

### v2

@@ 1 Introduction
Power side channels are well studied for classical computers [EPMS23, PR13, WSRW21, ABB+21, BDM+20, ABP19, WSW19, BCHC18, BYT17, BRN+13, EW14, MOPT12]. There are even platforms [OC14] for analyzing power side channels. However, understanding power side channels for quantum computers has not been explored yet, which this work aims to address. We show how a malicious attacker can reconstruct a secret quantum circuit that is being executed, by simply measuring the power consumption.
@@ 1.1 Power Side-Channel Threats to Quantum Computers
The vulnerabilities in quantum computer controllers encompass more than just gate recovery. An attacker might discern the number of qubits used in a quantum program, a significant concern for algorithms like Quantum Approximate Optimization Algorithm (QAOA) [CK19]. Variational quantum algorithms [CAB+21] are notably sensitive to qubit count and circuit depth. Extracting such hyperparameters can reveal crucial information about the algorithm. Moreover, given the current state of quantum computing where inputs are hard-coded, an attacker could potentially extract sensitive input data from the circuit.
@@ 1.2 Lessons from Historical Technological Threats
The significance of vulnerabilities in quantum computer controllers cannot be overstated, considering the ongoing evolution of technological advancements and their inherent risks. For perspective, speculative execution attacks in classical computers were not recognized as threats until 2018 [LSG+20, KHF+20], despite having been operational since the commercialization of the technology by IBM and Intel in the 1990s. This delay between innovation and the identification of vulnerabilities underscores the necessity for proactive security measures.
@@ 3.1 Assumptions of Attacker Measurement
A weaker attacker could collect a single total power trace over all channels, but not distinguish the power traces of each channel. Such attackers can attempt Circuit Reconstruction attack by collecting a single total power trace (see Figure 7). In particular, there is a trend to have multiple AWGs being part of same physical device. For example, in QICK [STW+22] framework, FPGAs are used for waveform generation and one FPGA can generate many control pulses. We believe that going forward total power side-channel attacks may be most realistic as attacker may not easily get power traces for individual channel generated by the FPGA, but can easily measure total power consumption of the FPGA, and thus get total power trace of all the channels. Nevertheless, we explore both per-channel and total power side-channel attacks to understand their potential threats.
@@ 5.1 Single-shot Per-Channel Power Side-Channel Attack
where $v_{i}(x)$ and $Total_{A_{P}}(x)_{i}$ are the $i$-th elements of $v(x)$ and $Total_{A_{P}}(x)$ respectively. However, we evaluated various distances and metrics on randomly generated circuits including Euclidean distance and found that the Jensen-Shannon distance is the most suitable one for quantifying the candidate instructions against the measurements. This distance metric is commonly used to measure the dissimilarity between probability distributions and is well-suited for the task of comparing the probability distributions of pulse sequences. It has been applied to genome comparison[SJWK09, IHS10] in protein surface comparison [OR03], in machine learning [GPAM+20] and particularly in the analysis of the similarity between two quantum states [OBL22]. To compute the Jensen-Shannon distance, we first convert the trace and pulse data of a candidate instruction to two probability vectors $P$ and $Q$ by normalizing the amplitudes to turn them into discrete probability distributions. We then calculate the distance between the two probability vectors by computing the Jensen-Shannon distance (metric) between two probability arrays. This is the square root of the Jensen-Shannon divergence ($\sqrt{{\rm JSD}}$[ES03].
@@
In per-channel single trace attacks, we use the Kullback-Leibler divergence (KLD) and the Jensen-Shannon divergence (JSD) and its distance ($\sqrt{\text{JSD}}$) to distinguish two probability distributions. Therefore, here we provide a brief introduction to divergences, distances and metrics. For more details, we refer the reader to [DDDD09, HIK+14]. A metric $d$ on a set $\chi$ is a function $d:\chi\times\chi\rightarrow\mathbb{R}_{\geq 0}$ such that for any $x,y,z\in\chi$ the following properties are satisfied:
$d(x,y)$ $\geq 0$ (non-negativity)
$d(x,y)$ $=0\text{ if and only if }x=y$ (identity of indiscernibles)
$d(x,y)$ $=d(y,x)$ (symmetry)
$d(x,y)$ $\leq d(x,z)+d(z,y)$ (triangle inequality)